a car travels 2 3/4 hours at an average speed of 60 km/hour and then for 2 1/4 hours at an average speed of 80km/hour. whats is the average speed of the whole journey?
i have worked out:
2 3/4 hours= 2 hours, 45 mins
2 1/4 hours= 2 hours and 15 mins

Answer by mananth(16949)

a car travels 2 3/4 hours at an average speed of 60 km/hour and then for 2 1/4 hours at an average speed of 80km/hour. whats is the average speed of the whole journey?

2 3/4 hours 60 km/hr
Distance traveled at this speed = 11/4 *60 = 165 km
2 1/4 hour 80 km/hr
Distance traveled at this speed = 9/4 * 80 = 180 miles
Total miles driven = 165+180= 345 km
Total time taken = 5 hours
Average speed for the total journey = total distance / total time = 345 / 5
69 km / hour
Ananth